# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

### v1

@@ I Introduction
Building on these ideas, we conclude this study by considering a real-world example. Throughout this paper, we emphasize, where possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics. By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e. a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales (as for the climate and weather system). Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [40]. For an early instance of case (ii), see the work of Charney (1948) [43], Charney et al. (1950) [44] and the discussion in Dalmedico (2001) [41]. The example we consider simulates the joint evolution of the sea surface temperature (SST) field and the net radiative flux at the top of the atmosphere (TOA). Our motivation lies in understanding the spatial structure of climate feedbacks by addressing the so-called pattern effect, i.e. the dependence of Earth’s energy imbalance on the spatial patterns of surface warming [30]. To this end, we develop a simplified stochastic neural model based on 20 SST modes and a global mean, scalar time series for the radiative flux. Unobserved degrees of freedom are parametrized through multiplicative (state-dependent) noise. We demonstrate that the model captures both the stationary statistics and the radiative fluxes response to imposed temperature perturbations in two climate change experiments. Such model can then be used to explore the linkages between surface temperature warming patterns and radiative fluxes, as well as a “reconstruction” tool, by predicting the change in one variable given prescribed changes in the other. Finally, we apply ideas from response theory to infer direct causal relationships between surface temperature and TOA radiative flux. These results show qualitative agreement with perturbation experiments using atmospheric-only climate models and align with current physical understanding of radiative feedbacks in the climate system.
@@ III.3.2 Stationary and perturbed statistics of the scalar model
These results indicate that the inclusion of multiplicative noise improves the reduced-order model’s ability to represent both stationary statistics and variance responses to perturbations. In the full triad system, a perturbation to $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$. Consequently, $x_{2}$ and $x_{3}$ feedback on $x_{1}$ at the consequent time step. This leads to a cascade of state-dependent interactions until the system relaxes back to stationarity. Such complex, state-dependent feedbacks cannot be represented by additive noise alone and require the inclusion of state dependent noise $\sigma(x)$, effectively parametrizing such interactions. However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right. The example illustrates that accurately capturing stationary statistics does not guarantee faithful reproduction of perturbed statistics.
@@
The presence of slowly decaying responses in the ensemble mean and variance to perturbations is indicative of a system that is no longer Markovian. The reduced-order stochastic scalar model is trained solely on the time series of $x_{1}(t)$. However, the time series $x_{1}(t)$ is generated by the full triad system in Eq. (5) and implicitly reflects the influence of the unobserved variables $x_{2}$ and $x_{3}$ on it. The variability of the unobserved variables $x_{2}$ and $x_{3}$ manifest as memory effects in the reduced order scalar model for $x_{1}$. One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [6] and later generalized by Kondrashov, Checkroun and Ghil (2014) [7]. The multilevel regression strategy method augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics, and can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2]. Ultimately, the remaining residuals become white in time, allowing the system to be closed using a Wiener process. Similarly, it could be possible to address the issue through recurrent neural networks [103]. An alternative, practical approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window. This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the the dynamics. Due to its simplicity and ease of implementation, this method is commonly used in the literature (see for example [4]) and is also adopted in the present study.
@@
We therefore consider again the time series $x_{1}(t)$ from the long simulation of the triad system in Eq. (5). As an additional step, we average the time series over a short temporal window of 10 time steps (recall that $dt=0.01$, so this is equivalent to $0.1$ Model Time Units). Importantly, this is not equivalent to sub-sampling the original dynamics by taking every 10th data point. Instead, averaging over a 10-step window smooths out fast variability and effectively defines a new, coarse-grained “more Markovian” system. We then re-train the neural models with additive and multiplicative noise on the new coarse-grained data. Next, we simulate the neural models over long timescales to analyze their stationary distributions, as well as their ensemble mean and variance responses to impulse perturbations. Results are shown in Figure 5. As before, the stochastic model with additive noise fails to capture the correct stationary statistics (Figure 5(a)). While the response in the ensemble mean improves relative to the non-coarse-grained case, the response in ensemble variance remains suppressed. In contrast, the model with multiplicative noise not only captures the stationary distribution and further improves the mean response, but now also closely reproduces the response in variance, capturing the peak magnitude and deviating only slightly in its temporal decay. Note that the stationary response in ensemble mean and variance to impulse perturbations of the original degree of freedom $x_{1}$ are the same in both Figures 4 and 5, the only thing changing in the latter is that the reduced-order model is defined in a coarse-grained dynamics.
@@ III.3.3 Main takeaways
Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector. This leads to two major challenges: (i) the resulting data-driven model will depend heavily on the specific variable chosen for modeling, and (ii) once a suitable slow variable is identified, the cumulative effect of the unobserved degrees of freedom must be appropriately parametrized. These issues, demonstrated at the “practical level” within the context of a low-dimensional model, become even more severe in realistic high-dimensional systems, which often lack a clear time-scale separation. These challenges are intrinsic to data-driven approaches for dynamical systems, and simply extending time series length or training increasingly large neural networks is unlikely to resolve them.
The core difficulty is conceptual: there is no rigorous mathematical framework for selecting the proper variables to model in real-world, high-dimensional dynamical systems. This poses fundamental limitations for purely data-driven approaches. In practice, such choices rely on physical intuition and prior knowledge. When possible, selecting the right variables amounts to identifying an appropriate coarse-grained description of the dynamics, including relevant fields and suitable spatial and temporal scales. Consequently, effective data-driven modeling of real-world complex systems hinges on deep physical insight and strong domain expertise.
@@ IV.2.1 Data, preprocessing and coarse-graining
The focus in on a state-of-the-art coupled climate model GFDL-CM4 [112]. The model offers data from a long control run simulating the coupled climate system for 600 years with constant $\text{CO}_{2}$ fixed at pre-industrial level. The ocean component is the MOM6 ocean model [113] with a horizontal grid spacing of 0.25 and 75 vertical layers. The atmospheric component is the AM4 model [114, 115] with a horizontal grid spacing of roughly 100 km and 33 vertical layers. The model considered in Eq. (10) will be trained on the long control run. The variables considered are two: SST and the global mean net radiative flux at the top of the atmosphere (hereafter, often referred to as TOA). The focus on global mean net radiative flux, rather than its spatiotemporal representation is motivated by the fact that such variable does not show large scale patterns of variability. It is in fact a fast variable in respect to SST and the main signal we are interested in is its global average. The SST field is considered only in the latitudinal range $[-45,45]$, a choice informed by earlier studies highlighting the prominence of radiative feedbacks in the tropics and subtropics [30, 108].
@@ IV.2.2 Stationary statistics and forced responses
Given the $n+1$ state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined in the preprocessing step above, we can now train the emulator defined in Eq. (10) on the long control run of the GFDL-CM4 model. The trained emulator is then used to simulate a 600 years long climate trajectory (same length as the one used in the original model GFDL-CM4). The stationary probability distributions of the first SST mode and of the net radiative flux at the TOA are found to be qualitatively well captured, see Figure 6(a,b). The largest discrepancies between the emulator and the original model are found for the autocorrelation functions. Autocorrelation functions are shown for the first SST mode (largely corresponding to the El Niño Southern Oscillation (ENSO) [119]) and for the TOA variable.
@@ IV.2.4 Main takeaways
In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature (SST) field and the global mean net radiative flux at the top of the atmosphere (TOA) in the GFDL-CM4 climate model. Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales. These simplifications led to a stable emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments. The model reconstruct the net TOA radiative flux when forced with a pre-computed SST trajectory. Beyond reconstruction, we illustrated that the model can be used to perform perturbation experiment, allowing us to enhance mechanistic understanding of the underlying physical processes. Findings are in qualitative agreement with results from traditional perturbation experiments in numerical models, giving additional confidence in the emulator’s ability to capture key interactions across variables.
@@ V Conclusions
In the first part of this paper, we addressed, at a pedagogical level, foundational issues in neural emulation of dynamical systems with climate applications in mind. We considered two scenarios: (i) when the full state vector is observed, and (ii) the more realistic case of partial observations. Our findings can be summarized as follows:
@@
In the second part of the paper, we built on these insights by constructing a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net top-of-atmosphere (TOA) radiative flux. The model was trained on long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies. We showed that the emulator reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in forced simulations when provided with forced SST trajectories. To deepen this analysis, we applied linear response theory to extract the direct causal links between temperature modes and TOA fluxes, targeting the so-called “pattern effect” [30, 29]. These links, when projected back onto the high-dimensional state space, were found to be consistent with recent studies. This suggests that the emulator’s skill in reconstructing TOA fluxes under forcing arises for the right physical reasons. In this way, focusing on a few dominant modes, on a limited range of time scales, and using stochastic parametrizations allowed us to build a skillful and interpretable emulator.
@@
Modeling real-world dynamical systems from data alone is severely constrained by high dimensionality and the fact that we typically observe only a subset of relevant variables [46, 32, 40]. This leads to the fundamental challenge of identifying, when possible, the proper variables to model, a problem recognized since the work of Onsager and Machlup (1953) [68] and highlighted in more recent studies [47, 48]. These challenges are especially pronounced in climate dynamics, where observations are always partial, and the underlying system is both high-dimensional and multiscale. In this context, the multiscale structure of the climate system can be an advantage, allowing us to model coarse-grained dynamics while treating unresolved processes stochastically. The need to formulate models tackling the effective dynamics of slow variables, rather than the full set of primitive equations, has a long history in climate science, even in non-data-driven settings where all governing equations are known. The early work of Charney and von Neumann, for example, demonstrated the conceptual and practical benefits of filtering out “meteorologically insignificant noise” and emphasized the need for interpretable and computationally efficient effective equations in metereology [123, 63, 43, 44]. These insights remain broadly relevant when developing neural emulators, especially in future attempts of building atmosphere-ocean coupled emulators. In purely data-driven settings, focusing on stochastic models that target coarse-grained processes offers a promising way forward by reconciling rigorous approaches such as reduced-order modeling with the flexibility and power of neural networks. Possible strategies include learning stochastic parameterizations directly from data using deep learning (as in this paper), guiding variable selection through causal inference [124, 125], incorporating information theory and response theory into the training phase [56], studying the evolution of the climate system in its “latent” dynamics [65] (i.e. on the attractor) and incorporating empirical knowledge and physics constraints in the learning phase [11, 7, 126, 25]. Finally, the rather hard difficulty (at times irreducible) of data-driven modeling for the coupled atmosphere–ocean system can at times be ameliorated by building models tailored to solve specific problems, by building on prior literature and insights, and developed with a clear understanding of the relevant variables and the spatiotemporal scales at play.

### v2

@@ I Introduction
Building on these ideas, we conclude this study by considering a real-world example. Throughout this paper, we emphasize, where possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics. By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e. a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales. Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [40]. For an early instance of case (ii), see the work of Charney (1948) [43], Charney et al. (1950) [44] and the discussion in Dalmedico (2001) [41]. The example we consider simulates the joint evolution of the sea surface temperature (SST) field and the net radiative flux at the top of the atmosphere (TOA). Our motivation lies in understanding the spatial structure of climate feedbacks by addressing the so-called pattern effect, i.e. the dependence of Earth’s energy imbalance on the spatial patterns of surface warming [30]. To this end, we develop a simplified stochastic neural model based on 20 SST modes and a global mean, scalar time series for the radiative flux. Unobserved degrees of freedom are parametrized through multiplicative (state-dependent) noise. We demonstrate that the model captures both the stationary statistics and the radiative fluxes response to imposed temperature perturbations in two climate change experiments. Such model can then be used to explore the linkages between surface temperature warming patterns and radiative fluxes, as well as a “reconstruction” tool, by predicting the change in one variable given prescribed changes in the other. Finally, we apply ideas from response theory to infer direct causal relationships between surface temperature and TOA radiative flux. These results show qualitative agreement with perturbation experiments using atmospheric-only climate models and align with current physical understanding of radiative feedbacks in the climate system.
@@ III.3.2 Stationary and perturbed statistics of the scalar model
These results indicate that the inclusion of multiplicative noise improves the reduced-order model’s ability to represent both stationary statistics and variance responses to perturbations. In the full triad system, a perturbation to $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$. Consequently, $x_{2}$ and $x_{3}$ feedback on $x_{1}$ at the consequent time step. This leads to a cascade of state-dependent interactions until the system relaxes back to stationarity. Such complex, state-dependent feedbacks cannot be represented by additive noise alone and require the inclusion of state dependent noise $\sigma(x)$, effectively parametrizing such interactions. However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right. We emphasize two key points:
  1. i.
    The analysis in Figure 4 highlights the importance of evaluating a model’s ability in predicting responses in higher-order moments, such as the variance. In the simple example discussed here, focusing solely on the ensemble mean response would lead to the incorrect conclusion that the two models respond to perturbations in the same way (see panels (b) and (e) in Figure 4).
  2. ii.
    Matching the stationary distribution quantitatively does not guarantee equally accurate responses to perturbations, as shown in Figure 4(d,e,f).
@@
The presence of slowly decaying responses in the ensemble mean and variance to perturbations is indicative of a system that is no longer Markovian. The reduced-order stochastic scalar model is trained solely on the time series of $x_{1}(t)$. However, the time series $x_{1}(t)$ is generated by the full triad system in Eq. (5) and implicitly reflects the influence of the unobserved variables $x_{2}$ and $x_{3}$ on it. The variability of the unobserved variables $x_{2}$ and $x_{3}$ manifest as memory effects in the reduced order scalar model for $x_{1}$. One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [6] and later generalized by Kondrashov, Checkroun and Ghil (2014) [7]. The multilevel regression strategy method augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics, and can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2]. Ultimately, the remaining residuals become white in time, allowing the system to be closed using a Wiener process. An alternative, practical approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window. This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the dynamics. Due to its simplicity and practicality, this method is commonly used in the literature (see e.g. [4]) and is also adopted in the present study.
@@
We therefore consider again the time series $x_{1}(t)$ from the long simulation of the triad system in Eq. (5). As an additional step, we average the time series over a short temporal window of 10 time steps (recall that $dt=0.01$, so this is equivalent to $0.1$ Model Time Units). Importantly, this is not equivalent to sub-sampling the original dynamics by taking every 10th data point. Instead, averaging over a 10-step window smooths out fast variability and effectively defines a new, coarse-grained “more Markovian” system. We then re-train the neural models with additive and multiplicative noise on the new coarse-grained data. Next, we simulate the neural models over long timescales to analyze their stationary distributions, as well as their ensemble mean and variance responses to impulse perturbations. Results are shown in Figure 5. As before, the stochastic model with additive noise fails to capture the correct stationary statistics (Figure 5(a)). While the response in the ensemble mean improves relative to the non-coarse-grained case, the response in ensemble variance remains suppressed. In contrast, the model with multiplicative noise not only captures the stationary distribution and further improves the mean response, but now also closely reproduces the response in variance, capturing the peak magnitude and deviating only slightly in its temporal decay. Note that the response in ensemble mean and variance to impulse perturbations of the original degree of freedom $x_{1}$ are the same in both Figures 4 and 5, the only thing changing in the latter is that the reduced-order model is defined in a coarse-grained dynamics.
@@ III.3.3 Main takeaways
Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector. This leads to two major challenges: (i) the resulting data-driven model will depend heavily on the specific variable chosen for modeling, and (ii) once a suitable slow variable is identified, the cumulative effect of the unobserved degrees of freedom must be appropriately parametrized. These issues, demonstrated at the “practical level” within the context of a low-dimensional model, become more severe in realistic high-dimensional systems, which often lack a clear time-scale separation. These challenges are intrinsic to data-driven approaches for dynamical systems, and simply extending time series length or training increasingly large neural networks is unlikely to resolve them. Crucially, proposing more sophisticated models can be effective only if the proper variables are first identified.
The core difficulty is conceptual: there is no rigorous mathematical framework for selecting the proper variables to model from data in real-world, high-dimensional dynamical systems. In practice, such choices rely on physical insight and domain expertise. In a spatiotemporal system like climate, this involves identifying an appropriate coarse-grained description of the dynamics, including relevant fields and a limited range of spatial and temporal scales. In the context of climate dynamics, it is sometimes possible (depending on the specific problem) to leverage the governing equations to guide variable selection. However, even in this case, one faces the practical necessity of resolving only a limited range of scales and introducing parametrizations. Within this framework, stochastic emulators offer a promising approach for data-driven simulation of climate systems and their sensitivity to external perturbations.
@@ IV.2.1 Data, preprocessing and coarse-graining
The focus is on a state-of-the-art coupled climate model GFDL-CM4 [111]. The model offers data from a long control run simulating the coupled climate system for 600 years with constant $\text{CO}_{2}$ fixed at pre-industrial level. The ocean component is the MOM6 ocean model [112] with a horizontal grid spacing of 0.25 and 75 vertical layers. The atmospheric component is the AM4 model [113, 114] with a horizontal grid spacing of roughly 100 km and 33 vertical layers. The model considered in Eq. (10) will be trained on the long control run. The variables considered are two: SST and the global mean net radiative flux at the top of the atmosphere (hereafter, often referred to as TOA). The focus on global mean net radiative flux, rather than its spatiotemporal representation is motivated by the fact that such variable does not show large scale patterns of variability. It is in fact a fast variable in respect to SST and the main signal we are interested in is its global average. The SST field is considered only in the latitudinal range $[-45,45]$, a choice informed by earlier studies highlighting the prominence of radiative feedbacks in the tropics and subtropics [30, 107].
@@ IV.2.2 Stationary statistics and forced responses
Given the $n+1$ dimensional state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined in the preprocessing step above, we can now train the emulator defined in Eq. (10) on the long control run of the GFDL-CM4 model. The trained emulator is then used to simulate a 600 years long climate trajectory (same length as the one used in the original model GFDL-CM4). The stationary probability distributions of the first SST mode and of the net radiative flux at the TOA are found to be qualitatively well captured, see Figure 6(a,b). The largest discrepancies between the emulator and the original model are found for the autocorrelation functions. Autocorrelation functions are shown for the first SST mode (largely corresponding to the El Niño Southern Oscillation (ENSO) [118]) and for the TOA variable.
@@ IV.2.4 Main takeaways
In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature (SST) field and the global mean net radiative flux at the top of the atmosphere (TOA) in the GFDL-CM4 climate model. Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales. These simplifications led to a stable and computationally efficient emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments. The model reconstruct the net TOA radiative flux when forced with a pre-computed SST trajectory. Beyond reconstruction, we illustrated that the model can be used to perform perturbation experiment, allowing us to enhance mechanistic understanding of the underlying physical processes. Findings are in qualitative agreement with results from traditional perturbation experiments in numerical models, giving additional confidence in the emulator’s ability to capture key interactions across variables.
@@ V Conclusions
In the first part of this paper, we addressed, at a pedagogical level, fundamental issues in neural emulation of dynamical systems with climate applications in mind. We considered two scenarios: (i) when the full state vector is observed, and (ii) the more realistic case of partial observations. Our findings can be summarized as follows:
@@
In addition, we argued that impulse response functions are a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations. This perspective enables us to move beyond evaluation metrics based solely on stationary statistics and instead consider the effect of external interventions. Assessing impulse responses is straightforward in the case of data-driven emulators built from simple numerical models, i.e. models that allow us to simulate thousands of ensemble members, as in the first part of this study. In such cases, one can actively perturb the reference model to generate a ground truth for the response operator. However, for models constructed from observational data, such direct perturbations are not possible. In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to estimate impulse response functions from data alone, at least at a qualitative level [38, 37, 35, 47, 62]. In particular, it has been shown that even simple approximations of FDT, such as the quasi-Gaussian approximation [37], often yield qualitatively reliable estimates of impulse responses in the ensemble mean, without the need to perturb the system directly; see, for example, [122, 123, 104]. That said, FDT itself is inferred from data and should therefore be viewed as a complementary, theory-based diagnostic (rather than a ground truth) allowing us to cross-validate the inferred responses of the emulator. In other words, consistency between the responses predicted by the emulator and those estimated via FDT can increase confidence in the model’s ability to capture the system’s response in the ensemble mean to external perturbations. In contrast, assessing impulse responses of higher-order moments (e.g., variance) via FDT remains substantially more challenging and calls for new methodological tools. A promising ongoing direction involves score-based generative modeling, as explored in [124, 125]. Another fruitful avenue lies in data-driven frameworks at the intersection of Markov State Modeling and Koopman operator theory [126, 127]. It is worth emphasizing that the central limitation discussed earlier, i.e. the need to choose the proper variables, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [47].
@@
Building on the theoretical insights introduced in the first part, the second part of the paper presents a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net top-of-atmosphere (TOA) radiative flux. The model was trained on long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies. We showed that the emulator reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in forced simulations when provided with forced SST trajectories. To deepen this analysis, we applied linear response theory to extract the direct causal links between temperature modes and TOA fluxes, targeting the so-called “pattern effect” [30, 29]. These links, when projected back onto the high-dimensional state space, were found to be consistent with recent studies. This suggests that the emulator’s skill in reconstructing TOA fluxes under forcing arises for the right physical reasons. In this way, focusing on a few dominant modes, on a limited range of time scales, and using stochastic parametrizations allowed us to build a skillful and interpretable emulator.
We emphasize that the fundamental limitations discussed here apply to data-driven models of complex dynamical systems that must both represent stationary statistics and respond faithfully to external perturbations. For Earth system models specifically, these considerations primarily concern climate emulators (particularly coupled atmosphere-ocean systems), and should not be conflated with the different challenge of weather forecasting using neural networks (e.g., [128]). Weather prediction focuses on short-term predictability where long-term stability and forced responses are of secondary importance.
@@
Modeling real-world dynamical systems from data alone is severely constrained by high dimensionality and the fact that we typically observe only a subset of relevant variables [46, 32, 40]. This leads to the fundamental challenge of identifying, when possible, the proper variables to model, a problem recognized since the work of Onsager and Machlup (1953) [68] and highlighted in more recent studies [47, 48]. These challenges are especially pronounced in climate dynamics, where observations are always partial, and the underlying system is both high-dimensional and multiscale. In this context, the multiscale structure of the climate system can be an advantage, allowing us to model coarse-grained dynamics while treating unresolved processes stochastically. The need to formulate models tackling the effective dynamics of slow variables, rather than the full set of primitive equations, has a long history in climate science, even in non-data-driven settings where all governing equations are known. The early work of Charney and von Neumann, for example, demonstrated the conceptual and practical benefits of filtering out “meteorologically insignificant noise” and emphasized the need for interpretable and computationally efficient effective equations in metereology [129, 63, 43, 44]. These insights remain conceptually relevant, especially the focus on effective dynamics, in the development of neural emulators for climate, including coupled atmosphere–ocean systems. In purely data-driven settings, focusing on stochastic models that target coarse-grained processes offers a promising way forward by reconciling rigorous approaches such as reduced-order modeling with the flexibility and power of neural networks. Possible strategies include learning stochastic parameterizations directly from data using deep learning (as in this paper), guiding variable selection through causal inference [130, 131], incorporating information theory and response theory into the training phase [56], studying the evolution of the climate system in its “latent” dynamics [65] (i.e. on the attractor) and incorporating empirical knowledge and physics constraints in the learning phase [11, 7, 132, 25]. Finally, the rather hard difficulty (at times irreducible) of data-driven modeling for the coupled atmosphere–ocean system can at times be ameliorated by building models tailored to solve specific problems, by building on prior literature and insights, and developed with a clear understanding of the relevant variables and the spatiotemporal scales at play.